OpenStudy (anonymous):

z-y+z=5 2y+3z=14 -3y+2z=5

OpenStudy (anonymous):

Find y and z?

OpenStudy (anonymous):

But how tho ?

OpenStudy (anonymous):

so as you look at the first equation, you see that there is z-y+z=5. which is the same as saying 2z-y=5. isolate y to itself, which becomes -y=-2z+5. and make it y=2z-5

OpenStudy (anonymous):

now you found y=2z-5, pick either the second or third equation to plug in y=2z-5 and solve for what z is, and simply plugging in what you get and then you can find what y and z equals.

OpenStudy (anonymous):

Thanks